OpenStudy (anonymous):

use the properties of logs to solve log_2(10x-x^2)=4.. the answers are in the back of the book which are 2 and 8.. but im not sure how to get there,, please post the steps! thank you!!

OpenStudy (gorv):

10x-x^2=2^4 10x-x^2=16 x^2-10x+16=0 x^2-8x-2x+16=0 x(x-8)-2(x-8)=0 (x-8)(x-2)=0 x=2 or x=8

OpenStudy (gorv):

@mathhelpplease9976

OpenStudy (anonymous):

how about the log_2 part i can just drop that?

OpenStudy (gorv):

2 is the base of log

OpenStudy (anonymous):

so i can just cross it out?

OpenStudy (gorv):

\[\log_{a} b=x \]

OpenStudy (gorv):

the b=a^x it is property of log

OpenStudy (gorv):

got it????

OpenStudy (anonymous):

sweet thank you! got it!!

OpenStudy (anonymous):

canu answer one more question

OpenStudy (gorv):

wlcme :)))

OpenStudy (gorv):

yeh sure

OpenStudy (anonymous):

solve log_2(log_3(x))=2 .. answer is 81

OpenStudy (gorv):

yeah buddy u grabbed it

OpenStudy (anonymous):

no no like that is the answer in the back of the book.. but i dont know how to work it out!! hahaha

OpenStudy (gorv):

log_3(x)=2^2=4 x=3^4=81

OpenStudy (anonymous):

where did u get 3^4

OpenStudy (anonymous):

got it thank you!!!!

OpenStudy (anonymous):

how about factor or use quadratic formula to solve: (5^x)^2-2(5^x)-1=0.. answers in the back of the book is log_5(1+sqrt(2)=ln(1+sqrt(2))/ln 5.. show work please

OpenStudy (gorv):

add and subtract 1 (5^x)^2-2(5^x)+1-1-1=0 (5^x)^2-2(5^x)+1-2=0 (5^x-1)^2=2 (5^x-1)=sqrt2 (5^x)=sqrt2+1 now take natural log both side ln(5^x)=ln(sqrt2+1) x*ln5=ln(sqrt2+1) x=ln(sqrt2+1)/ln5

OpenStudy (anonymous):

whyd you subtract and add 1 in the beginning

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

to make perfect square

OpenStudy (anonymous):

omg you are so msart.. how do you know all this math ?

OpenStudy (anonymous):

*smart

OpenStudy (anonymous):

how about.. solve the equation using the one to one property.. 3^x=27^x2.. answer is 0 and 1/3

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

are u asking the whole book lol

OpenStudy (gorv):

first practice on properties of log u will get everything

OpenStudy (anonymous):

hahaha ok. im just really bad at them and i have a test soon so trying to understand haha thanks

OpenStudy (anonymous):

how wouldi do that if there is x^2 thought on the right side like i know i should get common bases

OpenStudy (gorv):

3^x=27^x2 take log both side xlog3=x^2log27

OpenStudy (anonymous):

r u a math teacher or somethinggg haa

OpenStudy (gorv):

xlog3=x^2 * log3^3 ---------------------3^3=27 xlog3=3x^2*log3 x=3x^2 3x^2-x=0 x(3x-1)=0 x=0 or x=1/3

OpenStudy (gorv):

bro i just love maths....n i m doing my under grad

OpenStudy (anonymous):

in math?

OpenStudy (anonymous):

dude you are a life saver

OpenStudy (gorv):

no in engineering

OpenStudy (gorv):

i knoe that so i solved it man

OpenStudy (anonymous):

hahado u get paid for this or just for fun?